# Text-Visual Semantic Constrained AI-Generated Image Quality Assessment

### v2

@@ 5. Conclusion
Evaluating the quality of AGIs presents significant challenges, particularly in ensuring semantic alignment with textual prompts and accurately perceiving fine-grained visual details. This paper introduced SC-AGIQA, a unified framework designed to address these critical issues. By leveraging a Multimodal Large Language Model within the Text-assisted Semantic Alignment Module (TSAM) to generate a descriptive prompt for robust consistency checking, and incorporating Human Visual System priors through frequency domain analysis in the Frequency-domain Fine-grained Degradation Perception Module (FFDPM), our approach effectively tackles semantic misalignment and enhances sensitivity to subtle visual distortions. Extensive experiments conducted on the AGIQA-1k, AGIQA-3k, and AIGCIQA2023 benchmarks demonstrate the superiority of SC-AGIQA, establishing new state-of-the-art performance across all evaluated metrics. The results underscore the efficacy of synergizing advanced MLLM semantic understanding with perceptually-grounded visual feature extraction for comprehensive AGI quality assessment. While acknowledging the potential for refining descriptive prompt generation as discussed, SC-AGIQA offers a robust and effective solution, advancing the capabilities for reliable and nuanced evaluation of AI-generated image quality. We will further refine our approach by developing a more contextually-aware comparison conditioned on the original prompt.

### v3

@@ 5. Conclusion
Evaluating the quality of AGIs presents significant challenges, particularly in ensuring semantic alignment with textual prompts and accurately perceiving fine-grained visual details. This paper introduced SC-AGIQA, a unified framework designed to address these critical issues. By leveraging a Multimodal Large Language Model within the Text-assisted Semantic Alignment Module to generate a descriptive prompt for robust consistency checking, and incorporating Human Visual System priors through frequency domain analysis in the Frequency-domain Fine-grained Degradation Perception Module, our approach effectively tackles semantic misalignment and enhances sensitivity to subtle visual distortions. Extensive experiments conducted on three benchmarks demonstrate the superiority of SC-AGIQA, establishing new state-of-the-art performance across all evaluated metrics. The results underscore the efficacy of synergizing advanced MLLM semantic understanding with perceptually-grounded visual feature extraction for comprehensive AGI quality assessment. While acknowledging the potential for refining descriptive prompt generation as discussed, we will further enhance context-aware comparison conditioned on the original prompt.